# Quasi Zigzag Persistence: A Topological Framework for Analyzing Time-Varying Data

### v1

@@ Abstract
In this paper, we propose Quasi Zigzag Persistent Homology (QZPH) as a framework for analyzing time-varying data by integrating multiparameter persistence and zigzag persistence. To this end, we introduce a stable topological invariant that captures both static and dynamic features at different scales. We present an algorithm to compute this invariant efficiently. We show that it enhances the machine learning models when applied to tasks such as sleep-stage detection, demonstrating its effectiveness in capturing the evolving patterns in time-evolving datasets.
@@ 1 Introduction
Recently, topological Data Analysis (TDA) has emerged as a prominent field that can leverage topological information hidden behind the data into its analysis. Persistent Homology (PH), a cornerstone of TDA, provides a succinct method to extract and represent multiscale topological information. This capability has proved transformative in enhancing machine learning models, particularly where topological properties carry relevant information  Corbet et al. (2019); Chen et al. (2019); Carrière et al. (2020); Kim et al. (2020); Gabrielsson et al. (2020); Hofer et al. (2020); Zhao et al. (2020); Swenson et al. (2020); Carrière & Blumberg (2020); Vipond (2020); Bouritsas et al. (2022); Horn et al. (2022); Cang & Wei (2017); Zhang et al. (2022); Liu et al. (2022); Lesnick & Wright (2015). Time-varying data needs extensions of the classical one-parameter persistence homology along two directions. First, the temporal component of the data necessitates another parameter, thus bringing multi parameter persistence homology (MPH) into picture. Second, monotone filtrations usually associated with the standard PH computations cannot accommodate deletions that are required for processing time-varying data. This necessitates to use zigzag persistent homology (ZPH) which has been shown to capture dynamic topological features in time-series data Myers et al. (2023); Chen et al. (2021); Tinarrage et al. (2025). In effect, we need a combination of MPH and ZPH which allows PH (standard filtration) along one parameter and ZPH along the other. This requires changes to the structure of the underlying partially-ordered set (poset) to what we call quasi-zigzag poset resulting into Quasi Zigzag Persistent Homology (QZPH) introducing additional complexity and a new set of challenges.
While MPH captures richer topological information than PH, it presents a different set of challenges due to lack of a complete invariant. This has motivated TDA researchers to look for different invariants that, although not complete, capture sufficient topological information  Vipond (2020)Corbet et al. (2019); Carrière & Blumberg (2020); Scoccola et al. (2024); Loiseaux et al. (2023); Xin et al. (2023); Mukherjee et al. (2024). The juxtaposition of ZPH to MPH needs the zigzag generalizations of these methods.
@@ 2 Overview
$\mathbb{ZZ}\coloneqq\{(i,j)\colon i\in\mathbb{Z}^{\mathsf{op}},j\in\{i,i-1\}\}.$
@@
Next, we introduce the concepts of connectedness and intervals in posets which will be used for further exposition. We say a poset $P$ is connected if there is a sequence of points $p=r_{1},\ldots,r_{t}=q$ in $P$ so that $r_{i}\leq r_{i+1}$ or $r_{i+1}\leq r_{i}$ for every $i\in\{1,\ldots,t-1\}$. We say a subposet $I\subseteq P$ is convex if for every pair $p,q\in I$ and every $p\leq r\leq q$, we have $r$ in $I$. An interval $I$ in a poset $P$ is subposet which is both convex and connected. An interval $I$ with $m\geq 0$ minima and $n\geq 0$ maxima is called an $(m,n)$-interval. In the special case of $\mathbb{Z}^{2}$, every interval has a special form where the sets of minima and maxima along with other points constitute staircases that can be visualized as part of the boundary of the interval; see Figure 2.
@@
Let $M\colon P\to\mathbf{vec}$ be a persistence module from a finite and connected poset $P$. The restriction of $M$ to a subposet $I$ of $P$, $M|_{I}$, is the collection of vector spaces $M_{\mathbf{p}}$ for $\mathbf{p}\in I$ along with linear maps $M_{\mathbf{p}\to\mathbf{q}}$ for all comparable $\mathbf{p}\leq\mathbf{q}$ in $I$. Then, the generalized rank of $M$ over $I$ is defined as the rank of the canonical linear map from $\mathsf{lim}M|_{I}$ to $\mathsf{colim}M|_{I}$
@@ 3 Zz-Gril
Let $\mathbf{p}\in\mathbb{ZZ}\times\mathbb{Z}$ and $\delta\in\mathbb{Z}$ be given. Let $\boxed{\mathbf{p}}_{\delta}$ denote the $\delta$-square centered at $\mathbf{p}$, i.e., $\boxed{\mathbf{p}}_{\delta}\coloneqq\{\mathbf{z}\in\mathbb{ZZ}\times\mathbb{Z}\colon||\mathbf{p}-\mathbf{z}||_{\infty}\leq\delta\}$. Then, a worm centered at $\mathbf{p}$ with width $\delta$ is defined as the union of $\boxed{\mathbf{p}}_{\delta}$ with the two $\delta$-squares $\boxed{\mathbf{q}}_{\delta}$ centered at points $\mathbf{q}=\mathbf{p}\pm(\delta,-\delta)$ on the off-diagonal line segment. We denote the worm as $\boxed{\mathbf{p}}_{\delta}^{2}$. The superscript denotes the number of $\delta$-squares in the union apart from $\boxed{\mathbf{p}}_{\delta}$.
@@
Let $M$ be a quasi zigzag persistence module. Then, the ZigZag Generalized Rank Invariant Landscape is a function $\lambda^{M}\colon\mathbb{ZZ}\times\mathbb{Z}\times\mathbb{N}\to\mathbb{N}$ defined as
$\lambda^{M}(\mathbf{p},k)\coloneqq\sup\left\{\delta\geq 0\colon\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq k\right\},$
@@
Figure 2: The worm with blue boundary represents a worm centered at $\mathbf{p}$ with width $\delta=1$. The worm with green boundary represents an expanded worm, centered at $\mathbf{p}$ with width $\delta=2$.
@@ 3.1 Stability of Zz-Gril
Let $(\mathbb{ZZ}\times\mathbb{Z})$ denote the collection of all subposets in $\mathbb{ZZ}\times\mathbb{Z}$ such that their corresponding subposets in $\mathbb{Z}^{2}$ are intervals. For ϵ+italic-$\epsilon\in\mathbb{Z}_{+}$, the $\epsilon$-thickening of $I$ is defined as
$I^{\epsilon}\coloneqq\left\{\mathbf{r}\in\mathbb{ZZ}\times\mathbb{Z}\colon\exists\mathbf{q}\in I\text{ such that }||\mathbf{r}-\mathbf{q}||_{\infty}\leq\epsilon\right\}.$
@@
$\begin{split}d_{\mathcal{E}}^{\mathcal{L}}(M,N)\coloneqq\inf\limits_{\epsilon\geq 0}\{&\forall\boxed{\mathbf{p}}^{2}_{\delta}\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\\ &\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\text{ and }\\ &\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\}.\end{split}$
Note that $\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}$ contains the $\epsilon$-thickening of $\boxed{\mathbf{p}}^{2}_{\delta}$.
@@
We refer the reader to Appendix A for all the proofs.
@@ 4 Algorithm
We discuss the details of computing Zz-Gril in this section. We begin by proving that generalized rank over a subposet in $\mathbb{ZZ}\times\mathbb{Z}$ can also be computed by computing the number of full bars in the zigzag filtration along the “boundary” of the subposet. To be precise, we introduce the following concepts now. Let $P$ be any finite connected poset. We say a point $q\in P$ covers a point $p\in P$, denoted $p\prec q$ or $q\succ p$, if $p\leq q$ and there is no $r\in P$, $r\not=p,q$, so that $p\leq r\leq q$. A covering path $p\sim q$ between $p$ and $q$ is either a set of points $\{p=p_{0}\prec p_{1}\prec\cdots\prec p_{\ell}=q\}$ or $\{p=p_{0}\succ p_{1}\succ\cdots\succ p_{\ell}=q\}$. We define some special points in $P$. The set of minima in $P$ are the points $P_{min}=\{p\in I\ \,|\,\not\exists r\not=p\in P\mbox{ where }r\leq p\}$. Similarly, define the set of maxima $P_{max}$. For any two points $p,q\in P$, let $p\vee q$ and $p\wedge q$ denote their least upper bound (lub) and largest lower bound (llb) in $P$ if they exist.
@@ 5 Experiments
In this section, we report the results of testing Zz-Gril on various datasets. We begin by giving a detailed description about the experimental setup. Then, we give a brief description of the datasets, followed by the experimental results. We use the benchmark UEA multivariate time-series Bagnall et al. (2018) datasets to test Zz-Gril on multivariate time series data to show that Zz-Gril can be applied to datasets from various domains. Further, we test Zz-Gril on a targeted application of sleep-stage classification by performing experiments on ISRUC-S3 Khalighi et al. (2016) dataset. In all these experiments, we augment the topological information captured by Zz-Gril to one of the specifically tailored machine learning methods on the respective datasets and compare the performance. For each case, we select the machine learning model which has the highest performance to truly test and highlight the value of the topological information added by Zz-Gril to an already high-performing specifically tailored model.
@@ 5.1 Experimental Setup
Each data instance is a multivariate time-series. We convert multivariate time-series data into a quasi zigzag bi-filtration. The Zz-Gril framework takes in a quasi zigzag bi-filtration as input and provides a topological signature for the sequence as output. We augment the machine learning model with this topological information and train the model for classification. The framework is shown in Figure 6.
@@
Table 1: Accuracy and F-1 scores of augmenting Zz-Gril to STDP-GCN and testing on ISRUC-S3 sleep classification dataset. We report the overall accuracy, the overall F-1 score and sleep-stage-wise F-1 scores for each model.
Dataset/Methods ED-1NN DTW-1NN-I DTW-1NN-D MLSTM-FCN ShapeNet WEASEL+MUSE TapNet OS-CNN MOS-CNN TodyNet TodyNet+Zz-Gril
FingerMovements 0.550 0.520 0.530 0.580 0.589 0.490 0.530 0.568 0.568 0.570 0.660
Heartbeat 0.620 0.659 0.717 0.663 0.756 0.727 0.751 0.489 0.604 0.756 0.756
MotorImagery 0.510 0.390 0.500 0.510 0.610 0.500 0.590 0.535 0.515 0.640 0.660
NATOPS 0.860 0.850 0.883 0.889 0.883 0.870 0.939 0.968 0.951 0.972 0.961
SelfRegulationSCP2 0.483 0.533 0.539 0.472 0.578 0.460 0.550 0.532 0.510 0.550 0.600
Table 2: Acccuracy of augmenting Zz-Gril to TodyNet and testing on UEA Multivariate Time Series Classification Datasets.
5.2 UEA Multivariate Time Series Classification
The datasets are preprocessed and split into training and testing sets. We use these train-test splits and compare the performance, by augmenting Zz-Gril to TodyNet, Liu et al. (2024) with existing methods Li et al. (2021); Tang et al. (2022); Karim et al. (2019); Schäfer & Leser (2017). For this set of experiments, we convert the multivariate time series into a sequence of point clouds and compute Zz-Gril. We choose 5 datasets which have at least 7 multivariate time series. This ensures that each point cloud, in the sequence of point clouds, has at least 7 points, giving meaningful topological information. Refer to Figure 7 for a visualization of Zz-Gril on FingerMovements dataset processed as a sequence of point clouds. We report the results in Table 2. We can see that augmenting Zz-Gril improves the performance in most cases.
Dataset Time (Zz-Gril point clouds) Time (Zz-Gril graphs)
FingerMovements 569s 138s
NATOPS 413s 68s
SelfRegulationSCP2 153s 27s
Dataset TodyNet + Zz-Gril point clouds TodyNet + Zz-Gril graphs
FingerMovements 0.660 0.680
NATOPS 0.961 0.945
SelfRegulationSCP2 0.600 0.594
5.3 Sleep Stage Classification
We would like to clarify that the aim, for both sets of experiments, is primarily to show that an increase in accuracy signifies that Zz-Gril captures meaningful topological information which can be used to improve the existing models.
We report the computation times in Table 3. We can see from the table that Zz-Gril is, indeed, practical to use.

### v2

@@ Abstract
In this paper, we propose Quasi Zigzag Persistent Homology (QZPH) as a framework for analyzing time-varying data by integrating multiparameter persistence and zigzag persistence. To this end, we introduce a stable topological invariant that captures both static and dynamic features at different scales. We present an algorithm to compute this invariant efficiently. We show that it enhances the machine learning models when applied to tasks such as sleep-stage detection, demonstrating its effectiveness in capturing the evolving patterns in time-varying datasets.
@@ 1 Introduction
Recently, topological Data Analysis (TDA) has emerged as a prominent field that can leverage topological information hidden behind the data into its analysis. Persistent Homology (PH), a cornerstone of TDA, provides a succinct method to extract and represent multiscale topological information. This capability has proved transformative in enhancing machine learning models, particularly where topological properties carry relevant information  Corbet et al. (2019); Chen et al. (2019); Carrière et al. (2020); Kim et al. (2020); Gabrielsson et al. (2020); Hofer et al. (2020); Zhao et al. (2020); Swenson et al. (2020); Carrière & Blumberg (2020); Vipond (2020); Bouritsas et al. (2022); Horn et al. (2022); Cang & Wei (2017); Zhang et al. (2022); Bauer et al. (2023); Liu et al. (2022); Lesnick & Wright (2015). Time-varying data needs extensions of the classical one-parameter persistence homology along two directions. First, the temporal component of the data necessitates another parameter, thus bringing multi parameter persistence homology (MPH) Botnan & Lesnick (2023) into picture. Second, monotone filtrations usually associated with the standard PH computations cannot accommodate deletions that are required for processing time-varying data. This necessitates to use zigzag persistent homology (ZPH) Carlsson & de Silva (2010) which has been shown to capture dynamic topological features in time-series data Myers et al. (2023); Chen et al. (2021); Tinarrage et al. (2025); Hacquard & Lebovici (2024); Coskunuzer et al. (2024); Beltramo et al. (2022). In effect, we need a combination of MPH and ZPH which allows PH (standard filtration) along one parameter and ZPH along the other. This requires changes to the structure of the underlying partially-ordered set (poset) to what we call quasi-zigzag poset resulting into Quasi Zigzag Persistent Homology (QZPH) introducing additional complexity and a new set of challenges.
While MPH captures richer topological information than PH, it presents a different set of challenges due to lack of a complete invariant. This has motivated TDA researchers to look for different invariants that, although not complete, capture sufficient topological information  Vipond (2020)Corbet et al. (2019); Carrière & Blumberg (2020); Russold & Kerber (2024); Scoccola et al. (2024); Loiseaux et al. (2023); Xin et al. (2023); Mukherjee et al. (2024). The juxtaposition of ZPH to MPH needs the zigzag generalizations of these methods.
@@ 2 Overview
$\mathbb{ZZ}\coloneqq\{(i,j):i\in\mathbb{Z}^{\mathsf{op}},j\in\{i,i-1\}\}.$
@@
Next, we introduce the concepts of connectedness and intervals in posets which will be used for further exposition. We say a poset $P$ is connected if there is a sequence of points $p=r_{1},\ldots,r_{t}=q$ in $P$ so that $r_{i}\leq r_{i+1}$ or $r_{i+1}\leq r_{i}$ for every $i\in\{1,\ldots,t-1\}$. We say a subposet $I\subseteq P$ is convex if for every pair $p,q\in I$ and every $p\leq r\leq q$, we have $r$ in $I$. An interval $I$ in a poset $P$ is a subposet which is both convex and connected. An interval $I$ with $m\geq 0$ minima and $n\geq 0$ maxima is called an $(m,n)$-interval. In the special case of $\mathbb{Z}^{2}$, every interval has a special form where the sets of minima and maxima along with other points constitute staircases that can be visualized as part of the boundary of the interval; see Figure 2.
@@
Let $M:P\to\mathbf{vec}$ be a persistence module from a finite and connected poset $P$. The restriction of $M$ to a subposet $I$ of $P$, $M|_{I}$, is the collection of vector spaces $M_{\mathbf{p}}$ for $\mathbf{p}\in I$ along with linear maps $M_{\mathbf{p}\to\mathbf{q}}$ for all comparable $\mathbf{p}\leq\mathbf{q}$ in $I$. Then, the generalized rank of $M$ over $I$ is defined as the rank of the canonical linear map from $\mathsf{lim}M|_{I}$ to $\mathsf{colim}M|_{I}$
@@ 3 Zz-Gril
Let $\mathbf{p}\in\mathbb{ZZ}\times\mathbb{Z}$ and $\delta\in\mathbb{Z}$ be given. Let $\boxed{\mathbf{p}}_{\delta}$ denote the $\delta$-square centered at $\mathbf{p}$, i.e., $\boxed{\mathbf{p}}_{\delta}\coloneqq\{\mathbf{z}\in\mathbb{ZZ}\times\mathbb{Z}:||\mathbf{p}-\mathbf{z}||_{\infty}\leq\delta\}$. Then, a worm centered at $\mathbf{p}$ with width $\delta$ is defined as the union of $\boxed{\mathbf{p}}_{\delta}$ with the two $\delta$-squares $\boxed{\mathbf{q}}_{\delta}$ centered at points $\mathbf{q}=\mathbf{p}\pm(\delta,-\delta)$ on the off-diagonal line segment. We denote the worm as $\boxed{\mathbf{p}}_{\delta}^{2}$. The superscript denotes the number of $\delta$-squares in the union apart from $\boxed{\mathbf{p}}_{\delta}$.
@@
Let $M$ be a quasi zigzag persistence module. Then, the ZigZag Generalized Rank Invariant Landscape is a function $\lambda^{M}:\mathbb{ZZ}\times\mathbb{Z}\times\mathbb{N}\to\mathbb{N}$ defined as
$\lambda^{M}(\mathbf{p},k)\coloneqq\sup\left\{\delta\geq 0:\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq k\right\},$
@@
Figure 2: The worm with blue boundary represents a worm centered at $\mathbf{p}$ with width $\delta=1$. The worm with red boundary represents an expanded worm, centered at $\mathbf{p}$ with width $\delta=2$.
@@ 3.1 Stability of Zz-Gril
Let $(\mathbb{ZZ}\times\mathbb{Z})$ denote the collection of all subposets in $\mathbb{ZZ}\times\mathbb{Z}$ such that their corresponding subposets in $\mathbb{Z}^{2}$ are intervals.
@@
$d_{\mathcal{E}}^{\mathcal{L}}(M,N)\coloneqq\inf\left\{\epsilon\geq 0:\forall\boxed{\mathbf{p}}^{2}_{\delta}\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\text{ and }\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\right\}.$
@@
We refer the reader to Appendix A for the proofs of Proposition 3.4 and Theorem 3.5.
@@ 4 Algorithm
We discuss the details of computing Zz-Gril in this section. We begin by proving that generalized rank over a subposet in $\mathbb{ZZ}\times\mathbb{Z}$ can also be computed by computing the number of full bars in the zigzag filtration along the “boundary” of the subposet. To be precise, we introduce the following concepts now. Let $P$ be any finite connected poset. We say a point $q\in P$ covers a point $p\in P$, denoted $p\prec q$ or $q\succ p$, if $p\leq q$ and there is no $r\in P$, $r\not=p,q$, so that $p\leq r\leq q$. A covering path $p\sim q$ between $p$ and $q$ is either a set of points $\{p=p_{0}\prec p_{1}\prec\cdots\prec p_{\ell}=q\}$ or $\{p=p_{0}\succ p_{1}\succ\cdots\succ p_{\ell}=q\}$. We define some special points in $P$. The set of minima in $P$ are the points $P_{min}=\{p\in P\ \,|\,\not\exists r\not=p\in P\mbox{ where }r\leq p\}$. Similarly, define the set of maxima $P_{max}$. For any two points $p,q\in P$, let $p\vee q$ and $p\wedge q$ denote their least upper bound (lub) and largest lower bound (llb) in $P$ if they exist.
@@ 5 Experiments
In this section, we report the results of testing Zz-Gril on various datasets. We begin by giving a detailed description about the experimental setup. Then, we give a brief description of the datasets, followed by the experimental results. We use the benchmark UEA multivariate time-series Bagnall et al. (2018) datasets to test Zz-Gril on multivariate time series data to show that Zz-Gril can be applied to datasets from various domains. Further, we test Zz-Gril on a targeted application of sleep-stage classification by performing experiments on ISRUC-S3 Khalighi et al. (2016) dataset. In all these experiments, we augment the topological information captured by Zz-Gril to one of the specifically tailored machine learning methods on the respective datasets and compare.
@@ 5.1 Experimental Setup
Each data instance is a multivariate time-series which we convert into a quasi zigzag bi-filtration. The Zz-Gril framework takes in a quasi zigzag bi-filtration as input and provides a topological signature for the sequence as output. We augment the machine learning model with this topological information and train the model for classification. The framework is shown in Figure 6.
Dataset/Methods ED-1NN DTW-1NN-I DTW-1NN-D MLSTM-FCN ShapeNet WEASEL+MUSE OS-CNN MOS-CNN Zz-Gril
FingerMovements 0.550 0.520 0.530 0.580 0.589 0.490 0.568 0.568 0.590
Heartbeat 0.620 0.659 0.717 0.663 0.756 0.727 0.489 0.604 0.721
MotorImagery 0.510 0.390 0.500 0.510 0.610 0.500 0.535 0.515 0.580
NATOPS 0.860 0.850 0.883 0.889 0.883 0.870 0.968 0.951 0.850
SelfRegulationSCP2 0.483 0.533 0.539 0.472 0.578 0.460 0.532 0.510 0.522
Table 1: Acccuracy comparison of Zz-Gril with some existing methods on UEA Multivariate Time Series Classification Datasets. Bold entries denote the best model and gray the second best.
Dataset/Methods TapNet TapNet+Zz-Gril TodyNet TodyNet+Zz-Gril
FingerMovements 0.530 0.630 0.570 0.660
Heartbeat 0.751 0.751 0.756 0.756
MotorImagery 0.580 0.600 0.640 0.660
NATOPS 0.927 0.922 0.972 0.961
SelfRegulationSCP2 0.538 0.544 0.550 0.600
Table 2: Test acccuracy of augmenting Zz-Gril to TapNet and TodyNet on UEA Multivariate Time Series Classification Datasets.
5.2 UEA Multivariate Time Series Classification
The datasets are preprocessed and split into training and testing sets. For this set of experiments, we convert the multivariate time series into a sequence of point clouds and compute Zz-Gril. We choose 5 datasets which have at least 7 multivariate time series. This ensures that each point cloud, in the sequence of point clouds, has at least 7 points, giving meaningful topological information. Refer to Figure 7 for a visualization of Zz-Gril on FingerMovements dataset processed as a sequence of point clouds.
In Table 2, we compare the performance of augmenting Zz-Gril to two specifically tailored machine learning models for multivariate time series classification on UEA MTSC datasets, TapNet Zhang et al. (2020) and TodyNet Liu et al. (2024). We can see from the table that Zz-Gril is adding meaningful topological information on most datasets which is seen as an improvement in performance. We picked these two models for augmenting Zz-Gril because of two reasons: (i) availability and ease-of-use of codebase, (ii) these models already have good performance on UEA MTSC Datasets and we wanted to test the additional value of the topological information that Zz-Gril adds. In order to do a comprehensive evaluation of Zz-Gril framework, we compare Zz-Gril in isolation with other methods Li et al. (2021); Tang et al. (2022); Karim et al. (2019); Schäfer & Leser (2017) on UEA MTSC task. We report the results in Table 1. We see that Zz-Gril has comparable performance. Here, we would like to remind the reader that Zz-Gril is a general framework which does not need to be tailored for these datasets in particular. The aim for these experiments is to show that topological information is meaningful and is captured by Zz-Gril. Further, to show that we need both the spatial and the temporal information to be captured together, we compare Zz-Gril with standard zigzag persistence at various spatial scales. We report the results in Table 3. We can see that Zz-Gril performs better than standard zigzag persistence even across different scales.
Dataset Zigzag (scale = 0.5) Zigzag (scale = 0.7) Zizag (full graph) Zz-Gril
FingerMovements 0.490 0.490 0.530 0.590
Heartbeat - 0.682 0.687 0.721
MotorImagery 0.510 0.520 0.540 0.580
NATOPS - 0.805 0.816 0.850
SelfRegulationSCP2 0.483 0.489 0.500 0.522
Table 3: Accuracy of 1-parameter zigzag persistence at different scales compared with Zz-Gril. We use a Random Forest Classifier for classifying the topological signatures. The empty entries denote that the model does not train on the topological features because the features are not distinct enough at that scale.
We report the computation times in Table 4 in. We can see from the table that Zz-Gril is, indeed, practical to use.
Dataset Time (Zz-Gril point clouds) Time (Zz-Gril graphs)
FingerMovements 569s 138s
NATOPS 413s 68s
SelfRegulationSCP2 153s 27s
Dataset TodyNet + Zz-Gril point clouds TodyNet + Zz-Gril graphs
FingerMovements 0.660 0.680
NATOPS 0.961 0.945
SelfRegulationSCP2 0.600 0.594
5.3 Sleep Stage Classification
@@
Table 6: Accuracy and F-1 scores of augmenting Zz-Gril to STDP-GCN and testing on ISRUC-S3 sleep classification dataset. We report the accuracy and F-1 scores for each model.
We would like to clarify that the aim, for both sets of experiments, is primarily to show that an increase in accuracy upon augmentation signifies that Zz-Gril captures meaningful topological information which can be used to improve the existing models.